OpenStudy (gamerking1341):

What is the center and radius of the circle with equation (x-2)^2 + (y-3)^2=100?

OpenStudy (irishboy123):

Think about: \( X^2 + Y^2=10^2\) where: \( X = x-2, ~ Y = y-3\)